OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{\sqrt(n^{2}-1)}{n^{3}+2n^{2}+5}\] limit comparison test \[\frac{\sqrt(n^{2})}{n^{3}}\] \[\lim_{n \rightarrow\infty} \frac{1}{n} =0 \] convergent correct?

OpenStudy (blockcolder):

Yes.

OpenStudy (anonymous):

YESS!

OpenStudy (anonymous):

BEST ANSWER!

OpenStudy (anonymous):

thank you

OpenStudy (experimentx):

\[ \sum_{n=1}^{\infty} \frac{\sqrt(n^{2}-1)}{n^{3}+2n^{2}+5} \leq \sum_{n=1}^{\infty} \frac{\sqrt{n^{2}}}{n^{3}} = \sum_{n=1}^{\infty} \dfrac{1}{n^2}\] 1/n^2 converges ... don't compare it with 1/n ... 1/n diverges.